OpenStudy (anonymous):

Use the definition of subtraction to rewrite the subtraction equation as an addition equation. a) 34 - 6 = x b) x - 64 = 80

OpenStudy (anonymous):

help

OpenStudy (anonymous):

If 34 - 6 = x then what is x+6 ?

OpenStudy (anonymous):

I just need b I did a

OpenStudy (anonymous):

Ok well if x - 64 = 80 then what is x ?

OpenStudy (anonymous):

is it -64+x=80?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well maybe actually. I'm not entirely sure what the point of this exercise is.

OpenStudy (anonymous):

what is it I dont need it solved just need the equation as an addition prob.

OpenStudy (anonymous):

Are you supposed to be adding negatives

OpenStudy (anonymous):

Or are you supposed to use the equality fact.

OpenStudy (anonymous):

Im supposed to use the definiton of subtraction to make x - 64 = 80 as an additon prob....

OpenStudy (anonymous):

I can read. I just have no idea what that means.

OpenStudy (anonymous):

You can rewrite it as a lot of addition problems.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x + 0 = 144 x + 1 = 143 x + 2 = 142 .... x = 144 + 0 x = 143 + 1 x = 142 + 2 ... etc.

OpenStudy (anonymous):

Use the concept of addition to show that 8 > 2. wht about this one?

OpenStudy (anonymous):

And that's not even counting the options for doing both..

OpenStudy (anonymous):

That's pretty easy.. start by rewriting 8 as the sum of 2 and another number.

OpenStudy (anonymous):

Then you can see that it's greater.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

im geussing for b they want x+(-64)=80

OpenStudy (anonymous):

guessing*